OpenStudy (anonymous):

Subtract and simplify: (2p2q – 3pq2 + q3) – (p2q + q3) A. 2p2q – 3pq2 B. p2q – 3pq2 C. p2q – 2pq2 D. p2q – 3pq2+2q3 I thought this might be B but I wasn't sure...

OpenStudy (anonymous):

@imqwerty @mathstudent55

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@tootzrll @Pinkybottom67

OpenStudy (anonymous):

Someone....

OpenStudy (anonymous):

Anyone.....

OpenStudy (anonymous):

@Comrad

OpenStudy (mathstudent55):

When you subtract a polynomial from another polynomial, you drop the parentheses of the first one. Then to drop the parentheses of the second one, all signs inside the second parentheses change.

OpenStudy (anonymous):

Okay

OpenStudy (mathstudent55):

Here a simple example for you to see the change of signs. \((a + 2b) - (2a - 3b)\) \(=a + 2b - 2a + 3b\)

OpenStudy (anonymous):

Oh that makes sense.

OpenStudy (mathstudent55):

Notice carefully what was done from the first line tot eh second line above.

OpenStudy (mathstudent55):

The first polynomial remained exactly as it was, only the parentheses are no longer there. Notice (a + 2b) became a + 2b.

OpenStudy (anonymous):

Right, you just drop them and nothing changes.

OpenStudy (mathstudent55):

Now look at the second polynomial, the one being subtracted. To do the part -(2a - 3b), we drop the parentheses by changing every sign inside the parentheses. In other words, think of the minus sign to the left of the parentheses as being -1, and you distribute the -1 inside the parentheses.

OpenStudy (mathstudent55):

Inside the second parentheses, we had 2a - 3b. The 2a became -2a (the sign changed from 2 to -2). The -3b became + 3b ( the sign changed from -3 to +3).

OpenStudy (anonymous):

Oh okay

OpenStudy (mathstudent55):

Now that we have dealt with both sets of parentheses and have no parentheses left, we continue like we did for the sum of polynomials. We combine like terms.

OpenStudy (mathstudent55):

Here is where we left off above: \(=a + 2b - 2a + 3b\) Now we move like terms together: \(= a - 2a + 2b + 3b\) Finally, we combine like terms: \(= -a + 5b\)

OpenStudy (anonymous):

Oh I see.

OpenStudy (mathstudent55):

Now we can try your problem. Remember again the steps. Since the parentheses of the first polynomial are not necessary, we just drop the parentheses. To drop the parentheses of the polynomial we are subtracting, we change every sign inside the second parentheses. Then we combine like terms.

OpenStudy (anonymous):

1. 2p^2 - 3pq^2 + q^3 - p^2q + q^3

OpenStudy (mathstudent55):

Also, one more thing. To be like terms, two terms must have exactly the same variables and exactly the same exponents on those variables. Let's do your problem: \((2p^2q - 3pq^2 + q^3) - (p^2q + q^3)\)

OpenStudy (mathstudent55):

The first polynomial has parentheses which are not necessary. We simply drop them. \(=2p^2q - 3pq^2 + q^3 - (p^2q + q^3)\) Now we need to deal with the second parentheses. To drop the second parentheses, after the minus sign, we must change every sign inside the second parentheses. \(=2p^2q - 3pq^2 + q^3 - p^2q - q^3\)

OpenStudy (mathstudent55):

Notice the last \(+ q^3\) becomes \(- q^3\). Be careful with that.

OpenStudy (mathstudent55):

A negative outside parentheses changes ALL signs inside the parentheses. You see?

OpenStudy (anonymous):

Oh yeah

OpenStudy (mathstudent55):

We here here now: \(=2p^2q - 3pq^2 + q^3 - p^2q - q^3\) Now we move like terms together: \(= 2p^2q - p^2q -3pq^2 + q^3 - q^3 \) Ok so far?

OpenStudy (anonymous):

Yep!

OpenStudy (mathstudent55):

Now we combine like terms to finish our problem. \(= p^2q -3pq^2 \)

OpenStudy (mathstudent55):

You were correct again. B is the answer.

OpenStudy (mathstudent55):

Ok, sorry, but gtg.

OpenStudy (anonymous):

Oh great!! Thanks again so much, You helped a lot :)

OpenStudy (mathstudent55):

You're welcome.